# FFHFlow: A Flow-based Variational Approach for Learning Diverse Dexterous Grasps with Shape-Aware Introspection

### v2

@@ Abstract
Synthesizing diverse dexterous grasps from uncertain partial observation is an important yet challenging task for physically intelligent embodiments. Previous works on generative grasp synthesis fell short of precisely capturing the complex grasp distribution and reasoning about shape uncertainty in the unstructured and often partially perceived reality. In this work, we introduce a novel model that can generate diverse grasps for a multi-fingered hand while introspectively handling perceptual uncertainty and recognizing unknown object geometry to avoid performance degradation. Specifically, we devise a Deep Latent Variable Model (DLVM) based on Normalizing Flows (NFs), facilitating hierarchical and expressive latent representation for modeling versatile grasps. Our model design counteracts typical pitfalls of its popular alternative in generative grasping, i.e., conditional Variational Autoencoders (cVAEs) whose performance is limited by mode collapse and miss-specified prior issues. Moreover, the resultant feature hierarchy and the exact flow likelihood computation endow our model with shape-aware introspective capabilities, enabling it to quantify the shape uncertainty of partial point clouds and detect objects of novel geometry. We further achieve performance gain by fusing this information with a discriminative grasp evaluator, facilitating a novel hybrid way for grasp evaluation. Comprehensive simulated and real-world experiments show that the proposed idea gains superior performance and higher run-time efficiency against strong baselines, including diffusion models. We also demonstrate substantial benefits of greater diversity for grasping objects in clutter and a confined workspace in the real world.
I Introduction
@@
Performing diverse dexterous grasps on various objects is important to realize advanced human-like robotic manipulation. However, achieving such capability remains challenging due to the high dimensionality of the hand configuration space and the high variability in the object shape, not to mention the high perceptual uncertainty in the incomplete shape from partial observation. To address these challenges, we introduce a novel generative grasp synthesis model sketched in Figure 2.
Figure 1: An overview of the proposed idea: in this work, we propose a variational grasp sampler based on NFs, facilitating diverse dexterous grasps synthesis with shape-aware introspection.
@@
Previous efforts to address this problem have garnered significant interest in Conditional Variational Autoencoder(cVAE) [1, 2]. However, its performance is hampered by its fundamental issues of mode collapse  [3, 4, 5] and the often-used overly simple prior [6, 7]. These severely limit its ability to model the highly complex and inherently multi-modal grasp distribution. The lack of diversity significantly restricts the manipulability of robots operating in the common cluttered or confined space (e.g., a two-tier shelf in Figure 4) in daily life. Recent advances in the application of diffusion models [8] can potentially mitigate this issue but are struggling to achieve satisfactory run-time efficiency. More severely, scarce studies [9, 10] on generative dexterous grasping consider the necessity of having introspective capability against perceptual uncertainty raised by partial observation but with a run-time inefficient shape completion step. None of them takes into account the issue of Out-of-Distribution (OOD) objects with unknown object shapes uncovered by the training data, which easily causes unexpected consequences.
@@
In this work, we introduce a novel flow-based variational grasp synthesis model to overcome the aforementioned challenges. We first observed deficient generalization gain by simply using Conditional Normalizing Flows (cNF) to model the point cloud-conditioned grasp distribution due to suboptimal latent features. To address this, we introduce FFHFlow-lvm based on a novel flow-based Deep Latent Variable Models (DLVMs). In this model, two cNF are utilized, respectively, for an expressive input-dependent and meanwhile learnable prior [4, 7] (Prior Flow), and an elastic non-Gaussian likelihood function [3] (Grasp Flow) within the framework of DLVMs. With this design, we aim to conquer the over-regularization induced by the simple input-independent prior and the restricted isotropic Gaussian of the likelihood function in cVAE-based approaches.
To further facilitate shape-aware introspection, inspired by [11], we exploit the exact likelihoods of the Grasp Flow and Prior Flow to quantify the uncertainty for the incomplete shape of the partial point clouds, dubbed view-level shape uncertainty (Figure 6) and the unknown object shape of OOD objects (Figure 1) outside the training distribution, representing the object-level shape uncertainty. The information from Prior Flow enables our model to detect novel objects, preventing potential grasp failures. Moreover, we leverage Grasp Flow to assist grasp evaluation by striking a balance between increasing grasp quality captured by the grasp evaluator and lowering perceptual uncertainty due to partial observability.
@@
To summarize, we contribute with 1. a novel flow-based DLVMs that can address the limitations of its alternatives, such as cVAE and diffusion models; 2. an innovative grasp evaluation scheme that combines Grasp Flow and a grasp evaluator, and an OOD object detector based on Prior Flow; 3. a comprehensive experimental study of both in simulation and on the real robot to verify the proposed idea, i.e., object grasping from free space to clutter and confined space.
Figure 2: The flowchart of the proposed variational flow-based grasp sampler:FFHFlow-lvm  has a variational inference network and two cNF s for the prior and grasp generator. This design helps facilitate more expressive latent features under the framework of DLVMs (best viewed in color).
II Related Work
@@
Learning-based Grasp Synthesis. Deep generative models such as cVAE[1, 12, 2, 13], auto-regressive models [14], cNF[15, 16, 17], diffusion models [8, 18] and Generative Adversarial Networks (GANs)[19, 20] are widely adopted in grasp synthesis with two-jaw gripper and multi-fingered hands. Among them, some [2, 14] address the partial observation challenge by employing time-intensive shape completion modules, which significantly slow down inference, reducing their practical applicability. Other approaches [16, 17] assume the availability of complete object point clouds, a limitation in real-world scenarios. Similarly, diffusion models [8, 18] face runtime inefficiencies due to their iterative denoising processes, even though there are some techniques [21, 22] to accelerate it. In contrast, our method, similar to [15, 12], eliminates the need for resource-intensive shape completion models by encoding partial observation information directly into latent variables. This lightweight approach allows our model to effectively handle partial observations while maintaining efficiency and diversity in grasp synthesis. A diverse grasp synthesis enables real-world applications such as grasping in confined space, in clutter and dynamic grasping [23].
@@
Uncertainty-Aware Grasping. Prior research on uncertainty-aware grasping with deep learning (DL) is primarily focused on suction cups [24, 25] and parallel-jaw grippers [26, 27, 28]. Their motivations and applications stem from the characteristic of uncertainty estimation in DL [29], ranging from efficient adaptation [26] and exploration [24] to grasp evaluation [27, 25, 28]. Although with encouraging results, their investigation is limited to a rather simplistic setting, i.e., either using simplified end-effectors or restricting the analysis to 2D spaces. In contrast, non-DL based uncertainty-aware grasping considers shape analysis [30, 31, 32] and multi-fingered hands [33, 34], which exhibits an understudied gap. Our work aims to bridge this gap by developing a holistic learning-based grasping model that can retrospectively reason about shape-aware uncertainty against partially observed and unknown objects. It is worth noting that concurrent works [9, 10] share a similar spirit but rely on a time-consuming shape completion module, while our work is more real-time capable.
@@
III Preliminaries
Deep Latent Variable Models (DLVMs).
[…]
@@ III Preliminaries
Normalizing Flows.
NFs are known to be universal distribution approximators [39]. That is, they can model any complex target distribution $p^{*}(\mathbf{x})$ on a space $\mathcal{R}^{d}$ by defining $\mathbf{x}$ as a transformation $T_{\theta}:\mathcal{R}^{d}\rightarrow\mathcal{R}^{d}$ parameterized by $\theta$ from a well-defined base distribution $p_{u}(\textbf{u})$: $\mathbf{x}=T_{\theta}(\textbf{u})\text{~{}~{}where~{}~{}}\textbf{u}\sim p_{u}(\textbf{u})\,,$ where $\textbf{u}\in\mathcal{R}^{d}$ and $p_{u}$ is commonly chosen as a unit Gaussian. By designing $T_{\theta}$ to be a diffeomorphism, that is, a bijection where both $T_{\theta}$ and $T^{-1}_{\theta}$ are differentiable, we can compute the likelihood of the input $\mathbf{x}$ exactly based on the change-of-variables formula [48]: $p_{\theta}(\mathbf{x})=p_{u}(T_{\theta}^{-1}(\mathbf{x}))|\det(J_{T_{\theta}^{-1}}(\mathbf{x}))|$, where $J_{T_{\theta}^{-1}}(\mathbf{x})\in\mathcal{R}^{d\times d}$ is the Jacobian of the inverse $T_{\theta}^{-1}$ with respect to $\mathbf{x}$. The transformation $T_{\theta}$ can be constructed by composing a series of bijective maps denoted by $t_{i}$, $T_{\theta}=t_{1}\circ t_{2}\circ...\circ t_{n}$. When the target distribution is unknown, but samples thereof are available, we can estimate $\theta$ by minimizing the forward Kullback-Leibler Divergence (KLD), equivalent to minimizing the negative expected Log-Likelihood (LL):
@@
IV Flow-based Grasp Synthesis
IV-A Problem Formulation
This work aims to generate diverse grasps given a partial point cloud denoted by $\mathbf{x}\in\mathbb{R}^{N\times 3}$. A grasp configuration $\mathbf{g}\in\mathbb{R}^{d}$ is represented by the 15-DOF hand joint configuration $\boldsymbol{j}\in\mathbb{R}^{15}$ and the 6D palm pose $(\mathbf{R},\mathbf{t})\in SE(3)$. Given an empirical dataset of $N$ objects with $N_{i}$ corresponding possible grasps $\mathcal{D}=\{\mathbf{x}_{i},\{\mathbf{g}_{ik}\}_{k=1}^{N_{i}}\}_{i=1}^{N}$ drawn from the ground truth conditional distribution $p^{*}(\mathbf{g}|\mathbf{x})$, we train a probabilistic model $p_{\theta}(\mathbf{g}|\mathbf{x})$ parameterized by $\theta$ to approximate $p^{*}(\mathbf{g}|\mathbf{x})$. To facilitate shape-aware introspection [49], we exploit the log-likelihoods of $p_{\theta}(\mathbf{g}|\mathbf{x})$ to quantify the uncertainty caused by incomplete point clouds and unknown object geometries.
IV-B Flow-based Grasp Sampler: FFHFlow-cnf
A straightforward idea to learn the conditional distribution $p_{\theta}(\mathbf{g}|\mathbf{x})$ is directly employing the cNF [50] without defining hidden variables on the latent space shown in Figure 3a. To this end, we condition the flow transformation $T_{\theta}$ and the base distribution $p_{u}$ on the object point clouds $\mathbf{x}$, namely $T_{\theta|\mathbf{x}}:\mathcal{R}^{d}\times\mathcal{R}^{l}\rightarrow\mathcal{R}^{d}$ , where $l$ is the dimensionality of point cloud features and $d$ for the grasp representation. We encode each point cloud with a fixed Basis Point Set (BPS) according to [51], resulting in a feature vector $\mathbf{x_{b}}\in\mathbb{R}^{s}$ of fixed length $s$, before being fed into the feature extractor network $f_{\phi}(\mathbf{x_{b}}):\mathbb{R}^{s}\rightarrow\mathbb{R}^{l}$.
Limitation. Though FFHFlow-cnf achieved encouraging improvements in terms of diversity and accuracy when compared to the cVAE-based approach, we found it less generalizable with limited performance gain. We attribute the problem to the inadequate expressivity of the latent feature, especially when the model needs to understand the complicated relationships between the grasps and the partially observed point clouds of different objects. To address this problem, we introduce FFHFlow-lvm in the next sub-section, a flow-based variational sampler with a more expressive probabilistic representation in the latent space.
@@ IV-B Flow-based Grasp Sampler: FFHFlow-cnf
Figure 3: Graphical Models of Different Approaches. Shaded circles denote observed variables, and transparent ones for hidden variables. When compared to (a), DLVMs (b) and (c) have an additional hidden variable $\mathbf{z}$ for learning expressive latent features. However, the prior in the cVAE-based model (b) is an input-independent and simplistic Gaussian, while ours (c) is an input-dependent and more expressive cNF.
IV-C Flow-based Variational Grasp Sampler: FFHFlow-lvm
IV-C1 Learning Grasp Distribution via DLVMs
[…]
Variational Lower Bound

### v3

# FFHFlow: Diverse and Uncertainty-Aware Dexterous Grasp Generation via Flow Variational Inference

@@ Abstract
Synthesizing diverse, uncertainty-aware grasps for multi-fingered hands from partial observations remains a critical challenge in robot learning. Prior generative methods struggle to model the intricate grasp distribution of dexterous hands and often fail to reason about shape uncertainty inherent in partial point clouds, leading to unreliable or overly conservative grasps. We propose FFHFlow, a flow-based variational framework that generates diverse, robust multi-finger grasps while explicitly quantifying perceptual uncertainty in the partial point clouds. Our approach leverages a normalizing flow-based deep latent variable model to learn a hierarchical grasp manifold, overcoming the mode collapse and rigid prior limitations of conditional Variational Autoencoders (cVAEs). By exploiting the invertibility and exact likelihoods of flows, FFHFlow introspects shape uncertainty in partial observations and identifies novel object structures, enabling risk-aware grasp synthesis. To further enhance reliability, we integrate a discriminative grasp evaluator with the flow likelihoods, formulating an uncertainty-aware ranking strategy that prioritizes grasps robust to shape ambiguity. Extensive experiments in simulation and real-world setups demonstrate that FFHFlow outperforms state-of-the-art baselines (including diffusion models) in grasp diversity and success rate, while achieving run-time efficient sampling. We also showcase its practical value in cluttered and confined environments, where diversity-driven sampling excels by mitigating collisions (Project Page).
Keywords: Dexterous Grasping, Normalizing Flows, Uncertainty-Awareness
1 Introduction
@@
Performing diverse dexterous grasps on various objects is important to realize advanced human-like robotic manipulation. However, achieving such capability remains challenging due to the high dimensionality of the hand configuration space and the high variability in the object shape, not to mention the high perceptual uncertainty in the incomplete shape from the partial observation.
@@
Previous efforts to address this problem have garnered significant interest in Conditional Variational Autoencoder(cVAE) [1, 2]. However, its performance is hampered by the issues of mode collapse  [3, 4, 5] and the often-used overly simple prior [6, 7]. These severely limit its ability to model the highly complex and inherently multi-modal grasp distribution. The lack of diversity significantly restricts the manipulability of robots operating in the common cluttered or confined space (e.g., a two-tier shelf in Figure 4(a)) in daily life. Recent advances in the application of diffusion models [8] can potentially mitigate this issue but are struggling to achieve satisfactory run-time efficiency. More severely, only few prior studies [9, 10] considered the necessity of having introspective capabilities against the perceptual uncertainty raised by partial observation. Unfortunately, such solutions come with a often slow shape completion module,
Figure 1: Method overview: in this work, we propose a variational grasp sampler based on NFs, generating diverse dexterous grasps with shape-aware introspection.
not to mention the issue of Out-of-Distribution (OOD) objects with unknown object shapes uncovered by the training data, which easily cause unexpected consequences, e.g., grasp failure.
@@
In this work, we introduce a novel flow-based variational grasp generation model to overcome the aforementioned challenges (Figure 1). To this end, we first observed deficient generalization gain by simply using Conditional Normalizing Flows (cNF) to model the point cloud-conditioned grasp distribution due to suboptimal latent features. To address this, we devise FFHFlow-lvm based on a novel flow-based Deep Latent Variable Models (DLVMs) (Figure 2). In this model, we first replace the commonly used isotropic Gaussian with an expressive input-dependent NF prior [4, 7] (Prior Flow). Further, we construct an elastic non-Gaussian likelihood function [3] (Grasp Flow) with another NF within the framework of DLVMs. With this design, we aim to conquer the over-regularization induced by the simple input-independent prior and the restricted form of the likelihood function (e.g., isotropic Gaussian) in cVAE.
To facilitate shape-aware introspection, inspired by [11], we exploit the exact likelihoods of NFs to represent the perceptual uncertainty (lower the more uncertain) related to the generated grasps. Specifically, the likelihoods of Grasp Flow account for the incomplete shape due to partial observability, namely view uncertainty (illustrated in Figure 5(a)). It predicts lower values for grasps toward the invisible views. Meanwhile, Prior Flow assigns lower likelihoods for OOD objects with distinct shapes to the objects in the training data, representing the object uncertainty. More noteworthy, the object uncertainty can be applied to promote robustness by detecting objects with novel geometry, which can easily result in grasp failures. Moreover, we leverage the view uncertainty to assist grasp evaluation and propose a new evaluation strategy based on a discriminative grasp evaluator.
@@
To summarize, we contribute with (1) a novel flow-based DLVMs that can address the limitations of its alternatives, such as cVAE and diffusion models, (2) a new way to represent the perceptual uncertainties based on the flow likelihoods, leading to an uncertainty-aware grasp evaluation strategy and (3) a comprehensive experimental study both in simulation and on the real robot to verify the proposed idea, i.e., object grasping from free space to clutter and confined space.
2 Related Work
@@
Learning-based Grasp Synthesis. Deep generative models such as cVAE[1, 12, 2, 13], auto-regressive models [14], cNF[15, 16, 17], diffusion models [8, 18, 19] and Generative Adversarial Networks (GANs) [20, 21] are widely adopted in grasp synthesis with two-jaw gripper and multi-fingered hands. Among them, some [2, 14] address the partial observation challenge by employing time-intensive shape completion modules, which significantly slow down inference, reducing their practical applicability. Other approaches either assume the availability of complete object point clouds [16], a limitation in real-world scenarios, or only focus on a 2-jaw gripper [17]. Similarly, diffusion models [8, 18, 19] face runtime inefficiencies due to their iterative denoising processes, even though there are some techniques [22, 23] to accelerate it. In contrast, our method, similar to [15, 12], eliminates the need for resource-intensive shape completion models by encoding partial observation information directly into latent variables. This lightweight approach allows our model to effectively handle partial observations while maintaining efficiency and diversity in grasp synthesis. A diverse grasp synthesis enables real-world applications such as grasping in confined space, in clutter and dynamic grasping [24].
@@
Uncertainty-Aware Grasping. Prior research on uncertainty-aware grasping with deep learning (DL) is primarily focused on suction cups [25, 26] and parallel-jaw grippers [27, 28, 29]. Their motivations and applications stem from the characteristic of uncertainty estimation in DL [30], ranging from efficient adaptation [27] and exploration [25] to grasp evaluation [28, 26, 29, 31, 32]. Although with encouraging results, their investigation is limited to a rather simplistic setting, i.e., either using simplified end-effectors or restricting the analysis to 2D spaces. In contrast, non-DL based uncertainty-aware grasping considers shape analysis [33, 34, 35] and multi-fingered hands [36, 37], which exhibits an understudied gap. Our work aims to bridge this gap by developing a holistic learning-based grasping model that can retrospectively reason about shape-aware uncertainty against partially observed and unknown objects. Some concurrent works [9, 10] share a similar spirit but rely on a time-consuming shape completion module, while our work is more real-time capable.
@@
3 Preliminaries
Deep Latent Variable Models (DLVMs). In the context of modeling the unknown true data distribution $p^{*}(\mathbf{x})$ with a model $p_{\theta}(\mathbf{x})$ parameterized by $\theta$ based on a dataset $\mathcal{D}=\{\mathbf{x}_{i}\}^{N}_{i=1}$, latent variables $\mathbf{z}$ are usually introduced for discovering fine-grained factors controlling the data generating process or increasing the expressivity of the model $p_{\theta}(\mathbf{x})$. Latent variables $\{\mathbf{z}_{i}\}^{N}_{i=1}$ are part of the model but hidden and unobservable in the dataset. The resulting marginal probability is: $p_{\theta}({\mathbf{x}})=\int p_{\theta}(\mathbf{x}|\mathbf{z})p_{\theta}(\mathbf{z})d\mathbf{z}$. When $p_{\theta}(\mathbf{x},\mathbf{z})$ is parameterized by Deep Neural Networks (DNNs), we term the model DLVMs [49, 3]. The difficulty of learning such models with Maximum Likelihood Estimation (MLE) lies in the intractability of the integral in the marginal probability for not having an analytic solution or efficient estimator. To remedy this, Variational Inference (VI) [50] provides a tractable lower bound of the marginal likelihood $p_{\theta}(\mathbf{x})$ to optimize by approximating the real posterior $p_{\theta}(\mathbf{z}|\mathbf{x})$ with an approximate one $q_{\phi}(\mathbf{z}|\mathbf{x})$:
@@ 3 Preliminaries
Figure 2: Model Architecture: at training time, the approximate posterior of the latent variable is inferred while the derived variational lower bound is maximized. During inference, Prior flow generates latent samples conditioned on the input point clouds, based on which Grasp Flow predicts the final grasps (best viewed in color).
Normalizing Flows. NFs are known to be universal distribution approximators [42]. That is, they can model any complex target distribution $p^{*}(\mathbf{x})$ on a space $\mathcal{R}^{d}$ by defining $\mathbf{x}$ as a transformation $T_{\theta}:\mathcal{R}^{d}\rightarrow\mathcal{R}^{d}$ parameterized by $\theta$ from a well-defined base distribution $p_{u}(\textbf{u})$: $\mathbf{x}=T_{\theta}(\textbf{u})\text{~~where~~}\textbf{u}\sim p_{u}(\textbf{u})\,$, where $\textbf{u}\in\mathcal{R}^{d}$ and $p_{u}$ is commonly chosen as a unit Gaussian. By designing $T_{\theta}$ to be a diffeomorphism, that is, a bijection where both $T_{\theta}$ and $T^{-1}_{\theta}$ are differentiable, we can compute the likelihood of the input $\mathbf{x}$ exactly based on the change-of-variables formula [51]: $p_{\theta}(\mathbf{x})=p_{u}(T_{\theta}^{-1}(\mathbf{x}))|\det(J_{T_{\theta}^{-1}}(\mathbf{x}))|$, where $J_{T_{\theta}^{-1}}(\mathbf{x})\in\mathcal{R}^{d\times d}$ is the Jacobian of the inverse $T_{\theta}^{-1}$ with respect to $\mathbf{x}$. The transformation $T_{\theta}$ can be constructed by composing a series of bijective maps denoted by $t_{i}$, $T_{\theta}=t_{1}\circ t_{2}\circ...\circ t_{n}$. When the target distribution is unknown, but samples thereof are available, we can estimate $\theta$ by minimizing the forward Kullback-Leibler Divergence (KLD), equivalent to minimizing the negative expected Log-Likelihood (LL):
@@
4 Flow-based Grasp Synthesis
4.1 Problem Formulation
This work aims to generate diverse grasps given a partial point cloud denoted by $\mathbf{x}\in\mathbb{R}^{N\times 3}$. A grasp configuration $\mathbf{g}\in\mathbb{R}^{d}$ is represented by the 15-DOF hand joint configuration $\boldsymbol{j}\in\mathbb{R}^{15}$ and the 6D palm pose $(\mathbf{R},\mathbf{t})\in SE(3)$. Given an empirical dataset of $N$ objects with $N_{i}$ corresponding possible grasps $\mathcal{D}=\{\mathbf{x}_{i},\{\mathbf{g}_{ik}\}_{k=1}^{N_{i}}\}_{i=1}^{N}$ drawn from the ground truth conditional distribution $p^{*}(\mathbf{g}|\mathbf{x})$, we train a probabilistic model $p_{\theta}(\mathbf{g}|\mathbf{x})$ parameterized by $\theta$ to approximate $p^{*}(\mathbf{g}|\mathbf{x})$. To facilitate shape-aware introspection [52], we exploit the log-likelihoods of $p_{\theta}(\mathbf{g}|\mathbf{x})$ to represent the uncertainty caused by incomplete point clouds and unknown object shapes.
4.2 Flow-based Grasp Sampler: FFHFlow-cnf
A straightforward idea to learn the conditional distribution $p_{\theta}(\mathbf{g}|\mathbf{x})$ is directly employing the cNF [53] without defining hidden variables on the latent space shown in Figure 3(a). To this end, we condition the flow transformation $T_{\theta}$ and the base distribution $p_{u}$ on the object point clouds $\mathbf{x}$, namely $T_{\theta|\mathbf{x}}:\mathcal{R}^{d}\times\mathcal{R}^{l}\rightarrow\mathcal{R}^{d}$ , where $l$ is the dimensionality of point cloud features and $d$ for the grasp representation. We encode each point cloud with a fixed Basis Point Set (BPS) according to [54], resulting in a feature vector $\mathbf{x_{b}}\in\mathbb{R}^{s}$ of fixed length $s$, before being fed into the feature extractor network $f_{\phi}(\mathbf{x_{b}}):\mathbb{R}^{s}\rightarrow\mathbb{R}^{l}$.
Limitation: Though FFHFlow-cnf achieved encouraging improvements in terms of diversity and accuracy when compared to the cVAE-based approach, we found it less generalizable with limited performance gain. We attribute the problem to the inadequate expressivity of the latent feature (explained in Section 3.3 of the appendix), especially when the model needs to understand the complicated relationships between the grasps and the partially observed point clouds of different objects. To address this problem, we introduce FFHFlow-lvm in the next sub-section, a flow-based variational sampler with a more expressive probabilistic representation in the latent space.
4.3 Flow-based Variational Grasp Sampler: FFHFlow-lvm
@@
Figure 3: Graphical illustration. Shaded and unshaded circles denote observed and hidden variables, while arrows for dependencies. When compared to (a), DLVMs (b) and (c) have an extra hidden variable $\mathbf{z}$ for expressive latents learning. Though, the prior of $\mathbf{z}$ in cVAE (b) is an input-independent and simplistic Gaussian, while that of ours (c) is an input-dependent and more elastic cNF.
4.3.1 Learning Grasp Distribution via DLVMs
Our main idea is to introduce a latent variable into FFHFlow-cnf to increase the expressivity of the latent space. With the latent variable $\mathbf{z}$ shown in Figure 3(c), we have the following conditional likelihoods of the grasps $\mathbf{g}$ given a partially observed point cloud $\mathbf{x}$. It can be factorized: $p_{\theta}(\mathbf{g}|\mathbf{x})=\int p_{\theta}(\mathbf{g}|\mathbf{x},\mathbf{z})p_{\theta}(\mathbf{z}|\mathbf{x})d\mathbf{z}$. The likelihood is intractable due to the integral over the latent variables. We first derive a tractable lower bound for optimization (derivation in the appendix). Then, each component in the model is explained as follows how the pitfalls in cVAE are overcome. To note that $\theta$ denote the parameters for both flows and $\phi$ for the variational network.